Tohban Report 2016-08-17
From RHESSI Wiki
|Start Date:||10 August 2016|
|End Date:||17 August 2016|
|Next Tohban:||Juan Carlos Martinez Oliveros|
|List all reports|
Solar activity remained low this week, with comparable flares to last week. There are currently 5 named ARs on disk with 4 in the western hemisphere that will rotate off in the coming week. There is one AR in the eastern hemisphere and maybe one AR rotating on disk, but it doesn't have much complexity. The forecast calls for the chance of an M-class event with <10%.
How many GOES flares occurred?
Flares above B, C, M, X class were 24 5 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 5 2 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 4 1 0 0
There were RHESSI flares/GOES flares 15 / 22
over the time range 11-Aug-16 17-Aug-16
Decimation is normal/vigorous for the full-time period. The SSR is emptying before the end of the pass sets.
Temperatures continued to slowly rise on the Cold Tips this week. It was decided at a combined average of 117.5K action would be taken. In this week we did not reach that condition, though CT1 had reached 117.5. Temperature rise is slowing, as we had hoped with the day fraction changes. CT1 = 117.4, CT2 = 115.4, CPC ~ 56C
On Aug 16th (Tuesday) a special meeting was called to discuss the temperatures in more detail. Among discussion was that we don't have direct evidence that we are tapping, though exceeding the power guidelines by ~119%. It was decided to try and experiment where we lowered power and watched the temperatures. If the temperature stayed the same, then we could say that the additional power isn't giving us additional cooling, likely because of tapping. If we see the temperature increase, this would indicate that the additional power is going towards cooling and we likely aren't tapping. The following is Mark's email summarizing the commanding and experiment:
"We performed the experiment today, and saw a rapid rise in the cold tip temperature. The cold tip temperature increased about .7 K in one orbit (mostly in about 15 minutes). We then put the cryocooler back to 76.4 W and saw the cold tip temperature come back down again. This clearly demonstrates that we are still getting useful work from the increased cryocooler power, although I don't believe it completely rules out mild tapping.
At tomorrow's meeting, we will need to ponder what we want to do next. The cryocooler is currently at ~119% of recommended power and the cold plate is getting close to 140 K. The eclipse length will not be changing much, over the next couple of weeks."
Commands from today:
16-229-20:45:26 CP1: 140.4 K 16-229-20:45:26 CP2: 139.0 K 16-229-20:45:26 CT1: 117.0 K 16-229-20:45:26 CT2: 115.2 K 16-229-20:45:26 TST: 183.1 K 16-229-20:45:26 cp1 - ct1 = 23.19 K 16-229-20:45:26 Cryo Power: 74.2 W 16-229-20:45:26 Cryo Accel: 10.6 mG
16-229-20:44:42 /icryomain power=142 16-229-20:44:53 /icryobal amplitude=194 16-229-20:45:29 /icryomain power=140 16-229-20:45:34 /icryobal amplitude=192 16-229-20:45:50 /icryomain power=139 16-229-20:45:53 /icryobal amplitude=191
16-229-21:02:20 CP1: 140.4 K 16-229-21:02:20 CP2: 139.0 K 16-229-21:02:20 CT1: 117.8 K 16-229-21:02:20 CT2: 115.9 K 16-229-21:02:20 TST: 182.8 K 16-229-21:02:20 cp1 - ct1 = 22.63 K 16-229-21:02:20 Cryo Power: 71.4 W 16-229-21:02:20 Cryo Accel: 10.1 mG
16-229-22:23:39 /icryomain power=140 16-229-22:23:51 /icryobal amplitude=192 16-229-22:24:06 /icryomain power=142 16-229-22:24:12 /icryobal amplitude=194 16-229-22:24:16 start write_ssr_info 16-229-22:24:33 /icryomain power=144 16-229-22:24:40 /icryobal amplitude=196
16-230-00:02:59 CP1: 140.5 K 16-230-00:02:59 CP2: 139.0 K 16-230-00:02:59 CT1: 117.1 K 16-230-00:02:59 CT2: 115.2 K 16-230-00:02:59 TST: 182.8 K 16-230-00:02:59 cp1 - ct1 = 23.26 K 16-230-00:02:59 Cryo Power: 76.6 W 16-230-00:02:59 Cryo Accel: 11.6 mG
During the general meeting we discussed the experiment to raise the power. After talking with Goddard project folks we came up with the following plan.
First experiment: Lower the power by 1W and have the power increase by 1W in an ATS load. This is to prove that we will be able to lower power again if there is an issue with commanding on the spacecraft due to increased noise.
Second experiment: raise the power by 5W for an orbit and check the level of cooling. If we reach a 1/3rd degree level of cooling during an orbit, leave it at the elevated power for 24 hours. Then have a special meeting to discuss what to do next. During the remainder of the pass, ops will try to continue loading ATS loads to check if we will have issues getting the load up.
There are numerous VC3 data gaps in the record but all of them look to be replayed and filled in on browser.
As temperature increases, so does the noise on the detectors. Though the noise isn't too much in the front segments, it was enough for the shutters to be stuck in twice this week, necessitating increases in the shutter out threshold. The shutter in threshold remained unchanged for pileup concerns. We also changed the fast thresholds on D3 front, which had little effect on the deadtime.
Recommend that after doing the raise temperature experiment to raise the rear thresholds.
- First shutter threshold increase
2016/224 21:36 UTC. The thin out increased from 0x04 to 0x05 or from 1.57 to 1.96 % deadtime.
The new attenuator thresholds are as follows:
0.00 % Min A0 5.10 % Max A0 1.96 % Min A1 7.06 % Max A1 0.78 % Min A2 14.90 % Max A2 2.35 % Min A3 100.00 % Max A3
- Second shutter threshold increase and D3 front fast rate increase
16-228-19:35:20 start idib_chg_thrshld (3, FRONT, FAST, 0x38) 16-228-19:35:24 /IDPUDUMPTABL TABLE=DIBTBL3 16-228-19:35:24 /IDPUTABLE3 OFFSET=FRONTFASTDAC 16-228-19:35:37 /IDPULOAD VALUE=0x38
16-228-19:36:15 /IDPUDUMPTABL SCIENCE 16-228-19:36:27 /IDPUTABLED OFFSET=MINL1 16-228-19:37:23 /IDPULOAD VALUE=0x06
The new attenuator thresholds are as follows:
0.00 % Min A0 5.10 % Max A0 2.35 % Min A1 7.06 % Max A1 0.78 % Min A2 14.90 % Max A2 2.35 % Min A3 100.00 % Max A3
- D3 front fast rate and D8 rear fast rate threshold increases
16-228-22:50:35 start idib_chg_thrshld(3, FRONT, FAST, 0x40) 16-228-22:50:39 /IDPUDUMPTABL TABLE=DIBTBL3 16-228-22:50:39 /IDPUTABLE3 OFFSET=FRONTFASTDAC 16-228-22:50:53 /IDPULOAD VALUE=0x40
16-228-22:52:25 start idib_chg_thrshld(8, REAR, FAST, 0x68) 16-228-22:52:28 /IDPUDUMPTABL TABLE=DIBTBL8 16-228-22:52:28 /IDPUTABLE8 OFFSET=REARFASTDAC 16-228-22:52:34 /IDPULOAD VALUE=0x68
- Third shutter threshold increase
16-229-00:30:11 /IDPUDUMPTABL SCIENCE 16-229-00:30:30 /IDPUTABLED OFFSET=MINL1 16-229-00:30:51 /IDPULOAD VALUE=0x07
0.00 % Min A0 5.10 % Max A0 2.74 % Min A1 7.06 % Max A1 0.78 % Min A2 14.90 % Max A2 2.35 % Min A3 100.00 % Max A3
|Night time data (fronts)||full time|
|Night time data (rears)||full time|
|Require extra passes?||No|
|Requirement for moving pointer?||No|
|Detector problems?||Nominal in the new standard.|